40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Mar 6 01:19:19 2020, gautam, Update, IOO, Excess laser frequency noise MCF.pdfExcessFreqNoise.pdfALSnoise.pdf
    Reply  Fri Mar 6 16:33:11 2020, gautam, Update, IOO, Excess laser frequency noise 
Message ID: 15260     Entry time: Fri Mar 6 16:33:11 2020     In reply to: 15259
Author: gautam 
Type: Update 
Category: IOO 
Subject: Excess laser frequency noise 

I did some preliminary debugging of this, and have localized the problem to the output path (after MC slow) on the IMC Servo card. Basically, I monitored the spectrum of the ALS beat frequency fluctuations under a few different conditions: 

  • With the BNC to the NPRO PZT disconnected, there is no noise. So the laser and the FSS phase correction EOM (which the ALS beat pickoff sees) are not responsible.
  • With the input to the Koji summing box disconnected, still no excess - so the summing box + Thorlabs HV driver are not responsible.
  • With the TTFSS output connected to the summing box, but with the input switch to the TTFSS box disabled (isolating the on-PSL table parts of the FSS system), still no excess.
  • With the input to the TTFSS enabled, and the BNC output of the IMC Servo card disconnected at 1X2, still no excess.
  • Finally, when I connect the BNC cable, the excess starts to show up.

Toggling C1:IOO-MC_FASTSW, which supposedly isolates the post-MC slow (a.k.a. MCL) part of the servo, I see no difference. I am also reasonably confident this switch itself works, because I can break the IMC lock by toggling it. So pending a more detailed investigation, I am forced to conclude that the problem originates in the part of the IMC servo board after the MCL pickoff. Some cabling was removed at 1X2 on Tuesday between the times when there was no excess and when it showed up, but it's hard to imagine how this could have created this particular problem.

ELOG V3.1.3-